# An Information-Theoretic Framework for Receiver Quantization in Communication

### v2

@@ Abstract
We investigate information-theoretic limits and design of communication under receiver quantization. Unlike most existing studies that focus on low-resolution quantization, this work is more focused on the impact of weak nonlinear distortion due to resolution reduction from high to low. We consider a standard transceiver architecture, which includes independent and identically distributed (i.i.d.) complex Gaussian codebook at the transmitter, and a symmetric quantizer cascaded with a nearest neighbor decoder at the receiver. Employing the generalized mutual information (GMI), an achievable rate under general quantization rules is obtained in an analytical form, which shows that the rate loss due to quantization is $\log\left(1+\gamma\mathsf{SNR}\right)$, where $\mathsf{SNR}$ is the signal-to-noise ratio at the receiver front-end, and $\gamma$ is determined by thresholds and levels of the quantizer. Based on this result, the performance under uniform receiver quantization is analyzed comprehensively. We show that the front-end gain control, which determines the loading factor (normalized one-sided quantization range) of quantization, has an increasing impact on performance as the resolution decreases. In particular, we prove that the unique loading factor that minimizes the mean square error (MSE) of the uniform quantizer also maximizes the GMI, and the corresponding irreducible rate loss is given by $\log\left(1+\mathsf{mmse}\cdot\mathsf{SNR}\right)$, where $\mathsf{mmse}$ is the minimum MSE normalized by the variance of quantizer input, and it is equal to the minimum of $\gamma$. A geometrical interpretation for the optimal uniform quantization at the receiver is further established. Moreover, by asymptotic analysis, we characterize the impact of biased gain control, showing how small rate losses decay to zero and providing approximations for the achievable rate under large bias. From asymptotic expressions of the optimal loading factor and $\mathsf{mmse}$, approximations and several “per-bit rules” for performance are also provided. Finally we discuss more types of receiver quantization and show that the consistency between achievable rate maximization and MSE minimization does not hold in general.
@@ I Introduction
The analog-to-digital conversion (ADC), including sampling and quantization, is essential for any digital receiver. The power dissipation of state-of-the-art ADCs increases four times as the resolution increases by one bit, while every doubling of the sampling rate leads to a one-bit loss of resolution [2, 3]. The impact of ADC on the performance has received increasing attention along with the recent evolution of wireless communications, in which several challenges are faced, such as the increasing processing speed due to the utilization of larger bandwidth in mmWave and higher frequencies, the increasing scale of hardware due to the use of massive multiple-input-multiple-output (MIMO), and the critical need for low cost energy-efficient devices in emerging scenarios, e.g., massive machine-type communications (mMTC).
A majority of the studies on ADC at communication receivers have focused on the performance and design under low-resolution output quantization, and one-bit quantization has been of particular interest due to its negligible power dissipation and simplicity of implementation, even without requiring automatic gain control (AGC). In such studies the end-to-end channel is highly nonlinear, typically incurring a substantial performance loss, and necessitating a rethinking of the transceiver design.
On the other side, the transceiver architecture used in present wireless systems is built without considering the effect of output quantization. It is thus necessary to ask, under such conventional transceiver architecture, how much is the loss caused by output quantization with moderate to high resolution? In other words, if a small loss in achievable rate is acceptable, how fine need the quantization be? Analytical results on these problems appear to be lacking. Moreover, limited resolution of quantization leads to new problems, e.g., sensitivity of performance to the error of gain control, residual interference in multiuser systems, and so on. These largely unexplored problems prompt us to revisit the topic of receiver quantization in communication in this work.
@@ I-A Related Work on Quantization at Communication Receivers
We begin from the impact of output quantization in the (discrete-time) additive white Gaussian noise (AWGN) channel, which is a benchmark model in communication theory. The performance gain of using more output quantization levels in coded transmission (i.e., soft-decision decoding) was observed very early in [4] via an information theoretic approach. In the classic textbook of Wozencraft and Jacobs [References, Chap. 6.2], cutoff rate analysis showed that, for equiprobable uniformly spaced pulse amplitude modulation (PAM) input, the degradation due to output quantization is approximately $2$ dB when the alphabet size equals to the number of quantization levels, and the degradation vanishes when the quantization becomes increasingly fine. Particularly, in the low-signal-to-noise-ratio (low-SNR) limit, hard-decision decoding (one-bit quantization that observes the sign of output) leads to a power loss of $\pi/2$ (approximately $2$ dB) [4, 5]; see also [References, Chap. 2.11 and 3.4]. [^1] Interestingly, the loss can be fully recovered if we replace the hard-decision decoder (a.k.a. sign quantizer) by a carefully designed asymmetric one-bit quantizer (a.k.a. threshold quantizer) and employ asymmetric input [7]. For $K$-level output quantization, it was proved that a discrete input of at most $K+1$ mass points suffices to achieve the constrained capacity [8].
In vector (MIMO) channels, the quantization loss in achievable rate can be very small at low-to-moderate SNR even if 11~33-bit output quantization is used. This fact was shown in information theoretic studies [9, 10, 11] by numerical examples for full-rank channels with multiplexing gains 22~44. Although the constrained capacity is still unknown even in the one-bit case, recent information theoretic studies have provided various results on MIMO systems with coarse output quantization [12, 13, 14, 15, 16, 17, 18, 19], which show that proper transceiver design is critical in realizing efficient communication in such systems. In particular, for one-bit output quantization, it was shown that the high-SNR capacity grows linearly with the rank of channel [13], while the low-SNR asymptotic power loss of $\pi/2$ still exists in the case of vector channel [12].
@@ I-B Related Work in Quantization Theory
The rich theory of quantization was surveyed comprehensively in [42], in which two well established asymptotic theories were emphasized. The first is Shannon’s information theoretic approach (rate distortion theory [43]), which places quantization in the framework of lossy source coding and focuses on the high-dimension regime, thereby shedding light on vector quantization. The second is the asymptotic quantization theory, which sheds light on quantizer design in the high-resolution regime. The asymptotic quantization theory is more relevant to receiver quantization in communication which typically does not employ coding or vector quantization. Some classical results in asymptotic quantization theory, especially those for uniform scalar quantization, are reviewed here. Although not applicable directly to receiver quantization, they can be helpful in deriving new results (e.g., Corollary 10 in Sec. IV), closely related to our work (e.g., we will show the role of mean square error (MSE) in achievable rate analysis), and comparable to our findings (e.g., a remark in Sec. VII-A).
A basic result known in [44] and [45] (rigorously proved in [46]) states that, for a high-resolution uniform quantizer with step size $\ell$, the MSE can be approximated by $\ell^{2}/12$. This yields the “$6$-dB-per-bit rule” that each additional bit in resolution reduces the MSE by $6.02$ dB. The rule reflects the impact of step size that causes granular distortion; but it ignores overload distortion due to finite quantization range. The quantization range is often represented by a parameter called loading factor [47], which is defined as the one-sided width of quantization range normalized by the standard deviation of the input of the quantizer. The interplay between granular distortion and overload distortion determines how the optimal loading factor scales with the resolution. The scaling law has been characterized in the seminal work [48] for several types of input densities. Take the Gaussian source as example. In [48] it has been shown that, for the optimal $2K$-level uniform quantization that minimizes the MSE, 1) the loading factor scales like $2\sqrt{\ln(2K)}$ (cf. the conventional “four-sigma” rule of thumb [45, 49, 47]), and 2) the granular distortion dominates and the overload distortion is asymptotically negligible. Further properties of the uniform quantization have been analyzed in [50, 51, 52]. For quantization at communication receivers, we need parallel results to characterize the optimal loading factor, which is essential for the design of AGC. We note that, although the importance of AGC design in the presence of output quantization has been recognized for a long time [6, 30], it was only investigated by numerical results in several works; see, e.g., [9, 10, 53, 54, 22].
@@ I-C1 Problem
In this paper, we consider communications in the presence of output quantization, and focus on the effect of resolution reduction on the performance of a standard transceiver architecture designed without considering that effect. Rather than considering only low-resolution output quantization (11~33 bits) as in most existing studies, we consider the entire region of resolution, especially the transition from high resolution (typically 88~$12$ bits or more so that performance loss is negligible) to low resolution. Since in the transition only weak to moderate nonlinearity is introduced, it is natural to keep the transceiver architecture unchanged rather than rebuild it. The considered standard transceiver architecture includes independent and identically distributed (i.i.d.) complex Gaussian codebook at the transmitter, and a uniform output quantizer cascaded with a (weighted) nearest neighbor decoder at the receiver, where the loading factor of the quantizer can be adjusted by gain control. See Sec. II for details. We choose such an architecture in view of the following facts.
@@
  • The achievable rate of complex Gaussian input approximates that of regular high-order modulation schemes such as quadrature amplitude modulation (QAM). In the AWGN channel, the high-SNR gap between their achievable rates is 1.53 dB, which can be further reduced by constellation shaping [60].
@@ I-C3 Summary of Contribution
  • The uniform receiver quantization is further studied by asymptotic analysis. In Sec. V, the impact of biased gain control is characterized by asymptotic behaviors and approximations of $I_{\textrm{GMI}}(L)$. Approximations of $I_{\textrm{GMI}}(L)$ for overload region ($L<L^{*}$) and underload region ($L>L^{*}$) of the loading factor are proposed, respectively. Specifically, in the high-resolution regime, we characterize the loading loss in achievable rate, showing that i) the loading loss due to overload distortion decays exponentially as the loading factor increases, and ii) the loading loss due to granular distortion decays quadratically as the step size decreases. In Sec. VI, focusing on the optimal uniform quantization, we provide a new approximation of $L^{*}$, an approximation of $I_{\textrm{GMI}}(L^{*})$, and several per-bit rules for performance metrics such as saturation rate and irreducible rate loss;
@@ II-A A Standard Transceiver Architecture
Note (channel discretization): In practical systems, the described transceiver architecture also includes Nyquist-type pulse shaping at the transmitter and matched filtering combined with symbol-rate sampling at the receiver. It has a discrete-time memoryless model at symbol-level as (4), if the channel does not introduce memory. Note that oversampling (more accurately, sampling faster than the symbol rate) may improve achievable rate in the presence of output quantization, especially in the case of one-bit quantization (see [62, 63] and [64, 31, 65, 66]). However, such performance improvement relies on several conditions, including significant nonlinearity (leading to frequency dispersion that can be utilized by oversampling), non-standard transceiver architecture (like new waveform design at the transmitter [65, 66]), and sufficiently high SNR. Receivers based on oversampling are beyond the scope of this paper since our focus is the effect of low-to-moderate nonlinearity on a standard transceiver architecture.
@@ II-B An Achievable Rate Formula from Generalized Mutual Information
Maximizing the GMI by optimizing the scaling factor $a$ yields the following result [References, Appendix C], which provides a general approach for achievable rate analysis in the presence of transceiver nonlinearity with known transition probability.
@@
We note that, in Proposition 1, if $\mathrm{E}[Y]\neq 0$, then a pre-processing of the channel output as $Y\to\tilde{Y}=Y-\mathrm{E}[Y]$ improves the GMI (14). When $\mathrm{E}[Y]=0$ has been satisfied (e.g., the output $Y_{n}$ in (4) when $q(\cdot)$ is a symmetric quantizer), we have $\Delta=|\rho_{XY}|^{2}$, where
@@
and $\alpha$ is called the Bussgang gain. It is straightforward to verify that $\mathsf{SNR}_{\mathrm{e}}=\Delta/(1-\Delta)$, so that the rate (19) is equal to the GMI expression (14) in Proposition 1. In [40] this interpretation is extended to a more general setting, which includes memoryless channel state, imperfect CSIR, and vector (rather than scalar) channel output (but still assume i.i.d. Gaussian input). In brief, when the receiver linearly combines elements of channel output into a scalar for nearest neighbor decoding, the linear combiner that maximizes the GMI is exactly the linear minimum MSE (LMMSE) estimator of the channel input upon observing the channel output, and the maximum GMI is consistent with the achievable rate estimation based on Bussgang decomposition; see [References, Sec. IV-C] for details. [^5] In the description in [References, Sec. IV-C], it was incorrectly stated that the distortion is zero-mean. This may not hold for general nonlinear channels. But this does not affect the conclusion since in the effective SNR (cf. (20)), the distortion power is defined as the second-order moment, instead of the variance, of distortion. (Discussion with Yuhao Liu of Tsinghua University is acknowledged) Since the interpretation relies on i.i.d. Gaussian codebook, nearest neighbor decoding rule, and certain receiver processing, the following issues should be noted.
  • To achieve accurate rate estimation, the AQNM should be derived rigorously according to the Bussgang decomposition. In fact, this is not always satisfied in the literature. For example, in multiantenna systems, treating the residual noise in the AQNM as i.i.d. Gaussian vector (omitting the correlation among its elements) leads to inaccurate results, as pointed out in, e.g., [20], [41] (which includes an example of significant overestimate of achievable rate), and [29].

### v3

@@ Abstract
We investigate information-theoretic limits and design of communication under receiver quantization. Unlike most existing studies that focus on low-resolution quantization, this work is more focused on the impact of weak nonlinear distortion due to resolution reduction from high to low. We consider a standard transceiver architecture, which includes an independent and identically distributed (i.i.d.) complex Gaussian codebook at the transmitter, and a symmetric quantizer cascaded with a nearest neighbor decoder at the receiver. Employing the generalized mutual information (GMI), an achievable rate under general quantization rules is obtained in an analytical form, which shows that the rate loss due to quantization is $\log\left(1+\gamma\mathsf{SNR}\right)$, where $\mathsf{SNR}$ is the signal-to-noise ratio at the receiver front-end, and $\gamma$ is determined by thresholds and levels of the quantizer. Based on this result, the performance under uniform receiver quantization is analyzed comprehensively. We show that the front-end gain control, which determines the loading factor (normalized one-sided quantization range) of quantization, has an increasing impact on performance as the resolution decreases. In particular, we prove that the unique loading factor that minimizes the mean square error (MSE) of the uniform quantizer also maximizes the GMI, and the corresponding irreducible rate loss is given by $\log\left(1+\mathsf{mmse}\cdot\mathsf{SNR}\right)$, where $\mathsf{mmse}$ is the minimum MSE normalized by the variance of quantizer input, and it is equal to the minimum of $\gamma$. A geometrical interpretation for the optimal uniform quantization at the receiver is further established. Moreover, by asymptotic analysis, we characterize the impact of biased gain control, showing how small rate losses decay to zero and providing approximations for the achievable rate under large bias. From asymptotic expressions of the optimal loading factor and $\mathsf{mmse}$, approximations and several “per-bit rules” for performance are also provided. Finally we discuss more types of receiver quantization and show that the consistency between achievable rate maximization and MSE minimization does not hold in general.
@@ I Introduction
The analog-to-digital conversion (ADC), including sampling and quantization, is essential for any digital receiver. The power dissipation of state-of-the-art ADCs increases four times as the resolution increases by one bit, while every doubling of the sampling rate leads to a one-bit loss of resolution [2, 3]. The impact of the ADC on the performance has received increasing attention along with the recent evolution of wireless communications, in which several challenges are faced, such as the increasing processing speed due to the utilization of a larger bandwidth in the mmWave and higher frequencies, the increasing scale of hardware due to the use of massive multiple-input-multiple-output (MIMO), and the critical need for low cost energy-efficient devices in emerging scenarios, e.g., massive machine-type communications (mMTC).
A majority of the studies on the ADC at communication receivers have focused on the performance and design under low-resolution output quantization, and one-bit quantization has been of particular interest due to its negligible power dissipation and simplicity of implementation, even without requiring automatic gain control (AGC). In such studies the end-to-end channel is highly nonlinear, typically incurring a substantial performance loss, and necessitating a rethinking of the transceiver design.
On the other side, the transceiver architecture used in present wireless systems is built without considering the effect of output quantization. It is thus necessary to ask, under such a conventional transceiver architecture, how much is the loss caused by output quantization with moderate to high resolution? In other words, if a small loss in achievable rate is acceptable, how fine need the quantization be? Analytical results on these problems appear to be lacking. Moreover, limited resolution of quantization leads to new problems, e.g., sensitivity of performance to the error of gain control, residual interference in multiuser systems, and so on. These largely unexplored problems prompt us to revisit the topic of receiver quantization in communication in this work.
@@ I-A Related Work on Quantization at Communication Receivers
We begin from the impact of output quantization in the (discrete-time) additive white Gaussian noise (AWGN) channel, which is a benchmark model in communication theory. The performance gain of using more output quantization levels in coded transmission (i.e., soft-decision decoding) was observed very early in [4] via an information theoretic approach. In the classic textbook of Wozencraft and Jacobs [References, Chap. 6.2], a cutoff rate analysis showed that, for an equiprobable uniformly spaced pulse amplitude modulation (PAM) input, the degradation due to output quantization is approximately $2$ dB when the alphabet size equals to the number of quantization levels, and the degradation vanishes when the quantization becomes increasingly fine. Particularly, in the low-signal-to-noise-ratio (low-SNR) limit, hard-decision decoding (one-bit quantization that observes the sign of output) leads to a power loss of $\pi/2$ (approximately $2$ dB) [4, 5]; see also [References, Chap. 2.11 and 3.4]. [^1] Interestingly, the loss can be fully recovered if we replace the hard-decision decoder (a.k.a. sign quantizer) by a carefully designed asymmetric one-bit quantizer (a.k.a. threshold quantizer) and employ asymmetric input [7]. For $K$-level output quantization, it was proved that a discrete input of at most $K+1$ mass points suffices to achieve the constrained capacity [8].
In vector (MIMO) channels, the quantization loss in achievable rate can be very small at low-to-moderate SNR even if 11~33-bit output quantization is used. This fact was shown in information theoretic studies [9, 10, 11] by numerical examples for full-rank channels with multiplexing gains 22~44. Although the constrained capacity is still unknown even in the one-bit case, recent information theoretic studies have provided various results on MIMO systems with coarse output quantization [12, 13, 14, 15, 16, 17, 18, 19], which show that proper transceiver design is critical in realizing efficient communication in such systems. In particular, for one-bit output quantization, it was shown that the high-SNR capacity grows linearly with the rank of channel [13], while the low-SNR asymptotic power loss of $\pi/2$ still exists in the case of the vector channel [12].
@@ I-B Related Work in Quantization Theory
The rich theory of quantization was surveyed comprehensively in [42], in which two well-established asymptotic theories were emphasized. The first is Shannon’s information theoretic approach (rate distortion theory [43]), which places quantization in the framework of lossy source coding and focuses on the high-dimension regime, thereby shedding light on vector quantization. The second is the asymptotic quantization theory, which sheds light on quantizer design in the high-resolution regime. The asymptotic quantization theory is more relevant to receiver quantization in communication which typically does not employ coding or vector quantization. Some classical results in asymptotic quantization theory, especially those for uniform scalar quantization, are reviewed here. Although not applicable directly to receiver quantization, they can be helpful in deriving new results (e.g., Corollary 10 in Sec. IV), closely related to our work (e.g., we will show the role of the mean square error (MSE) in the achievable rate analysis), and comparable to our findings (see, e.g., the remark in Sec. VII-A).
A basic result known in [44] and [45] (rigorously proved in [46]) states that, for a high-resolution uniform quantizer with step size $\ell$, the MSE can be approximated by $\ell^{2}/12$. This yields the “$6$-dB-per-bit rule” that each additional bit in resolution reduces the MSE by $6.02$ dB. The rule reflects the impact of the step size that causes granular distortion; but it ignores overload distortion due to finite quantization range. The quantization range is often represented by a parameter called loading factor [47], which is defined as the one-sided width of quantization range normalized by the standard deviation of the input of the quantizer. The interplay between granular distortion and overload distortion dictates the optimal loading factor for a given resolution. The asymptotic relationship between this optimal loading factor and the resolution has been characterized in the seminal work [48] for several types of input densities. Take the Gaussian source as an example. In [48] it has been shown that, for the optimal $2K$-level uniform quantization that minimizes the MSE, 1) the loading factor scales like $2\sqrt{\ln(2K)}$ (cf. the conventional “four-sigma” rule of thumb [45, 49, 47]), and 2) the granular distortion dominates and the overload distortion is asymptotically negligible. Further properties of the uniform quantization have been analyzed in [50, 51, 52]. For quantization at communication receivers, we need parallel results to characterize the optimal loading factor, which is essential for the design of the AGC. We note that, although the importance of AGC design in the presence of output quantization has been recognized for a long time [6, 30], it was only investigated by numerical results in several works; see, e.g., [9, 10, 53, 54, 22].
@@ I-C1 Problem
In this paper, we consider communications in the presence of output quantization, and focus on the effect of resolution reduction on the performance of a standard transceiver architecture designed without considering that effect. Rather than considering only low-resolution output quantization (11~33 bits) as in most existing studies, we consider the entire region of resolution, especially the transition from high resolution (typically 88~$12$ bits or more so that performance loss is negligible) to low resolution. Since in the transition only weak to moderate nonlinearity is introduced, it is natural to keep the transceiver architecture unchanged rather than rebuild it. The considered standard transceiver architecture includes an independent and identically distributed (i.i.d.) complex Gaussian codebook at the transmitter, and a uniform output quantizer cascaded with a (weighted) nearest neighbor decoder at the receiver, where the loading factor of the quantizer can be adjusted by gain control. See Sec. II for details. We choose such an architecture in view of the following facts.
@@
  • The achievable rate of a complex Gaussian input approximates that of regular high-order modulation schemes such as quadrature amplitude modulation (QAM). In the AWGN channel, the high-SNR gap between their achievable rates is 1.53 dB, which can be further reduced by constellation shaping [60].
@@ I-C3 Summary of Contribution
  • The uniform receiver quantization is further studied by asymptotic analysis. In Sec. V, the impact of biased gain control is characterized by asymptotic behaviors and approximations of $I_{\textrm{GMI}}(L)$. Approximations of $I_{\textrm{GMI}}(L)$ for overload region ($L<L^{*}$) and underload region ($L>L^{*}$) of the loading factor are proposed, respectively. Specifically, in the high-resolution regime, we characterize the loading loss in achievable rate, showing that i) the loading loss due to the overload distortion decays exponentially as the loading factor increases, and ii) the loading loss due to the granular distortion decays quadratically as the step size decreases. In Sec. VI, focusing on the optimal uniform quantization, we provide a new approximation of $L^{*}$, an approximation of $I_{\textrm{GMI}}(L^{*})$, and several per-bit rules for performance metrics such as saturation rate and irreducible rate loss;
@@ II-A A Standard Transceiver Architecture
Note (channel discretization): In practical systems, the described transceiver architecture also includes Nyquist-type pulse shaping at the transmitter and matched filtering combined with symbol-rate sampling at the receiver. It has a discrete-time memoryless model at symbol-level as (4), if the channel does not introduce memory. Note that oversampling (more accurately, sampling faster than the symbol rate) may improve the achievable rate with output quantization, especially in the case of one-bit quantization (see [62, 63] and [64, 31, 65, 66]). However, such performance improvement relies on several conditions, including significant nonlinearity (leading to frequency dispersion that can be utilized by oversampling), non-standard transceiver architecture (like new waveform design at the transmitter [65, 66]), and sufficiently high SNR. Receivers based on oversampling are beyond the scope of this paper since our focus is the effect of low-to-moderate nonlinearity on a standard transceiver architecture.
@@ II-B An Achievable Rate Formula from Generalized Mutual Information
Maximizing the GMI by optimizing the scaling factor $a$ yields the following result [References, Appendix C], which provides a general approach for the achievable rate analysis under transceiver nonlinearity with known transition probability.
@@
We note that, in Proposition 1, if $\mathrm{E}[Y]\neq 0$, then a pre-processing of the channel output as $Y\to\tilde{Y}=Y-\mathrm{E}[Y]$ improves the GMI (14). When $\mathrm{E}[Y]=0$ has been satisfied (e.g., when $q(\cdot)$ is a symmetric quantizer, the output $Y_{n}$ in (4) satisfies $\mathrm{E}[Y_{n}]=0$), we have $\Delta=|\rho_{XY}|^{2}$, where
@@
and $\alpha$ is called the Bussgang gain. It is straightforward to verify that $\mathsf{SNR}_{\mathrm{e}}=\Delta/(1-\Delta)$, so that the rate (19) is equal to the GMI expression (14) in Proposition 1. In [40] this interpretation is extended to a more general setting, which includes a memoryless channel state, imperfect CSIR, and a vector (rather than scalar) channel output (but still assume i.i.d. Gaussian input). In brief, when the receiver linearly combines elements of the channel output into a scalar for nearest neighbor decoding, the linear combiner that maximizes the GMI is exactly the linear minimum MSE (LMMSE) estimator of the channel input upon observing the channel output, and the maximum GMI is consistent with the achievable rate estimation based on Bussgang decomposition; see [References, Sec. IV-C] for details. [^5] In the description in [References, Sec. IV-C], it was incorrectly stated that the distortion is zero-mean. This may not hold for general nonlinear channels. But this does not affect the conclusion since in the effective SNR (cf. (20)), the distortion power is defined as the second-order moment, instead of the variance, of distortion. (Discussion with Yuhao Liu of Tsinghua University is acknowledged) Since the interpretation relies on i.i.d. Gaussian codebook, the nearest neighbor decoding rule, and a certain receiver processing, the following issues should be noted.
  • To achieve accurate rate estimation, the AQNM should be derived rigorously according to the Bussgang decomposition. In fact, this is not always satisfied in the literature. For example, in multiantenna systems, treating the residual noise in the AQNM as i.i.d. Gaussian vector (omitting the correlation among its elements) leads to inaccurate results, as pointed out in, e.g., [20], [41] (which includes an example of significant overestimate of the achievable rate), and [29].